OpenStudy (anonymous):

Can some one help me with this problem please. (1/2)^3=

OpenStudy (anonymous):

Can you please type a Fraction

OpenStudy (kropot72):

\[(\frac{1}{2})^{3}=\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}=?\]

OpenStudy (anonymous):

= 1/2.Right

OpenStudy (anonymous):

= 1/8

OpenStudy (kropot72):

Yes. 1/8 is the correct answer.

ganeshie8 (ganeshie8):

\( (\frac{1}{2})^{3}=\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}= \frac{1\times1\times 1}{2\times 2\times 2} = ?\)

OpenStudy (anonymous):

1/8